# Mutual Information Density of Massive MIMO Systems over Rayleigh-Product Channels

### 05 Jan 2024 v4

@@ VI Conclusion and Future Works
This paper characterized the MID of massive MIMO systems over Rayleigh-product channels by setting up a CLT utilizing RMT. Based on the CLT, we performed the FBL analysis and analyzed the impact of rank-deficiency. Specifically, we first derived the upper and lower bounds for the packet error probability, and then obtained their high and low SNR approximations, which explicitly show the impact of rank-deficiency. The results in this work degenerate to the FBL results for Rayleigh channels when the number of scatterers approaches infinity, the IBL results for Rayleigh-product channels when the blocklength approaches infinity, and the IBL results for Rayleigh channels, when both the number of scatterers and the blocklength go to infinity. The derived results can be extended to double-scattering channels and IRS-aided MIMO channels with Rayleigh fading [48] by the approach in Appendix C with the aid of the computation results in [31]. Considering challenges in CSI acquisition for large-scale MIMO systems, it is also meaningful to extend the results to the case with imperfect CSI setting [49].
@@ Appendix B Mathematical Tools and Useful Results
1. Nash-Poincaré Inequality [19, Eq. (18)][35, Proposition 2.5]. Denote $\mathbb{x}=[x_{1},...,x_{N}]^{T}$ as a complex Gaussian random vector satisfying $\mathbb{E}\mathbb{x}=\mathbb{0}$, $\mathbb{E}\mathbb{x}\mathbb{x}^{T}=\mathbb{0}$, and $\mathbb{E}\mathbb{x}\mathbb{x}^{H}=\mathbb{\Omega}$. $f=f(\mathbb{x},\mathbb{x}^{*})$ is a $\mathbb{C}^{1}$ complex function such that both itself and its derivatives are polynomially bounded. Then, the variance of $f$ satisfies the following inequality,
@@
$\mathbb{E}[x_{i}f(\mathbb{x},\mathbb{x}^{*})]=\sum_{m=1}^{N}[\mathbb{\Omega}]_{i,m}\mathbb{E}[\frac{\partial f(\mathbb{x},\mathbb{x}^{*})}{\partial x_{m}^{*}}].$ (55)
@@
$\mathbb{E}[x_{i}f(\mathbb{x},\mathbb{x}^{*})]=\mathbb{E}[\frac{\partial f(\mathbb{x},\mathbb{x}^{*})}{\partial x_{m}^{*}}].$ (56)
@@
Given Assumption A and defining ${\mathbb{Q}}(z)=(z{\mathbb{I}}_{N}+{\mathbb{H}}{\mathbb{H}}^{H})^{-1}$, there holds true that for finite $x,z>0$, $m\geq 0$ and matrix ${\mathbb{B}}$,
$\mathrm{Var}(\frac{\operatorname{Tr}{\mathbb{B}}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}}{M})\leq\frac{K_{m}\operatorname{Tr}{\mathbb{B}}{\mathbb{B}}^{H}(1+2mz^{-2}+2x^{-2})}{M^{3}z^{2m}x^{2}},$ (57)
$\mathrm{Var}(\frac{\operatorname{Tr}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}{\mathbb{H}}{\mathbb{B}}{\mathbb{H}}^{H}}{M})\leq\frac{K_{m}\operatorname{Tr}{\mathbb{B}}{\mathbb{B}}^{H}(3+2x^{-2}+2mz^{-2})}{M^{3}z^{2m}x^{2}},$ (58)
$\mathrm{Var}(\frac{\operatorname{Tr}{\mathbb{Q}}(x){\mathbb{Q}}(z)^{m}{\mathbb{H}}{\mathbb{B}}{\mathbb{H}}^{H}}{M})\leq\frac{K_{m}\operatorname{Tr}{\mathbb{B}}{\mathbb{B}}^{H}(2+2x^{-2}+2mz^{-2})}{M^{3}z^{2m}x^{2}},$ (59)
$\mathrm{Var}(\frac{\operatorname{Tr}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{B}}{\mathbb{H}}^{H})^{m}}{M})\leq\frac{K_{m}\operatorname{Tr}({\mathbb{B}}{\mathbb{B}}^{H})^{m}(2m+1+2x^{-2}+2mz^{-2})}{M^{3}x^{2}z^{2m}},$ (60)
$\mathrm{Var}(\frac{(\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}))^{m}{\mathbb{Q}}(x){\mathbb{H}}{\mathbb{B}}{\mathbb{H}}^{H}}{M})\leq\frac{K_{m}\operatorname{Tr}({\mathbb{B}}{\mathbb{B}}^{H})(m+2+2x^{-2}+2mz^{-2})}{M^{3}x^{2}z^{2m}}.$ (61)
@@
$\mathrm{Var}(\frac{\operatorname{Tr}{\mathbb{B}}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}}{M})\leq\frac{\mathbb{E}}{M^{2}L}[\sum_{i,j}|[{\mathbb{Y}}{\mathbb{H}}^{H}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}{\mathbb{B}}{\mathbb{Q}}(x)]_{j,i}|^{2}+|[{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}{\mathbb{B}}{\mathbb{Q}}(x){\mathbb{H}}{\mathbb{Y}}^{H}]_{i,j}|^{2}$ (62)
@@
$+|[{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}{\mathbb{B}}{\mathbb{Q}}(x)]_{j,i}|^{2}+|[{\mathbb{Q}}(z)^{m}{\mathbb{B}}{\mathbb{Q}}(x){\mathbb{Z}}]_{i,j}|^{2}]$
$+\frac{\mathbb{E}}{M^{3}}[\sum_{i,j}|[{\mathbb{H}}^{H}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}{\mathbb{B}}{\mathbb{Q}}(x){\mathbb{Z}}]_{j,i}|^{2}+|[{\mathbb{Z}}^{H}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}{\mathbb{B}}{\mathbb{Q}}(x){\mathbb{H}}]_{i,j}|^{2}$
$+\sum_{k=1}^{m}|[{\mathbb{H}}^{H}{\mathbb{Q}}^{k}(z){\mathbb{B}}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m-k+1}{\mathbb{Z}}]_{j,i}|^{2}+|[{\mathbb{Z}}^{H}{\mathbb{Q}}^{k}(z){\mathbb{B}}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m-k+1}{\mathbb{H}}]_{i,j}|^{2}]$
$\leq\frac{1}{M^{3}}[2K_{1}z^{-2m}x^{-4}\operatorname{Tr}{\mathbb{B}}{\mathbb{B}}^{H}+2K_{1}mz^{-2(m+1)}x^{-2}\operatorname{Tr}{\mathbb{B}}{\mathbb{B}}^{H}+2\operatorname{Tr}{\mathbb{B}}{\mathbb{B}}^{H}x^{-2}z^{-2m}$
$+2K_{2}z^{-2m}x^{-4}\operatorname{Tr}{\mathbb{B}}{\mathbb{B}}^{H}+2K_{2}mz^{-2(m+1)}x^{-2}\operatorname{Tr}{\mathbb{B}}{\mathbb{B}}^{H}]$
$\leq\frac{K_{m}\operatorname{Tr}({\mathbb{B}}{\mathbb{B}}^{H})^{m}(2m+1+2x^{-2}+2mz^{-2})}{M^{3}x^{2}z^{2m}},$
@@
$\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]}{M}=\frac{\omega\delta_{z}}{\delta(1+\delta_{z}\overline{\omega}\overline{\omega}_{z})\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}),$ (67)
$\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})]}{M}=\frac{\overline{\omega}\overline{\omega}_{z}\omega_{z}\delta_{z}}{\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}),$ (68)
@@
$=\frac{\delta\overline{\omega}\omega_{z}\overline{\omega}_{z}}{1+\delta\overline{\omega}\overline{\omega}_{z}}+\frac{M\overline{\omega}\omega_{z}^{2}\overline{\omega}_{z}}{L\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})^{2}\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}),$
@@
$+\frac{M\omega_{z}}{L\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]}{M}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{M^{2}z}),$
@@
$\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}}{M}=\frac{1}{M}\sum_{i,j}\mathbb{E}Y_{j,i}^{*}[{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}]_{j,i}=\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}-\mathbb{E}\frac{\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}}{M}\frac{\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}}{M}$ (73)
$-\mathbb{E}\frac{\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}\frac{\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}}{M}=\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}-\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}}{M}\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}}{M}$
$-\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}}{M}+\varepsilon_{H,1}+\varepsilon_{H,2},$
where $\varepsilon_{H,1}=-\frac{1}{M^{2}}\mathrm{Cov}(\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H},\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})$ and $\varepsilon_{H,2}=-\frac{1}{M^{2}}\mathrm{Cov}(\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H},\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})$. By using the Cauchy-Schwarz inequality and the variance control in (57) and (59) of Lemma 2, $|\varepsilon_{H,1}|$ and $|\varepsilon_{H,2}|$ can be bounded by
$|\varepsilon_{H,1}|\leq\frac{1}{M^{2}}\mathrm{Var}^{\frac{1}{2}}(\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})\mathrm{Var}^{\frac{1}{2}}(\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})={\mathcal{O}}(\frac{\mathcal{P}(z^{-1})}{M^{2}z^{2}}),$ (74)
$|\varepsilon_{H,2}|\leq\frac{1}{M^{2}}\mathrm{Var}^{\frac{1}{2}}(\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})\mathrm{Var}^{\frac{1}{2}}(\operatorname{Tr}{\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})={\mathcal{O}}(\frac{\mathcal{P}(z^{-1})}{M^{2}z}),$
@@
$\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}}{M}=(1-\omega\overline{\omega})\frac{\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}-\frac{\omega_{z}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}}{M}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}})$ (75)
=$\overset{(b)}{=}\frac{\overline{\omega}\overline{\omega}_{z}\mathbb{E}\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}}{M}+{\mathcal{O}}(\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}),$

### 09 Apr 2024 v5

@@ VI Conclusion and Future Works
This paper characterized the MID of massive MIMO systems over Rayleigh-product channels by setting up a CLT utilizing RMT. Based on the CLT, we performed the FBL analysis and analyzed the impact of rank-deficiency. Specifically, we first derived the upper and lower bounds for the packet error probability, and then obtained their high and low SNR approximations, which explicitly show the impact of rank-deficiency. The results in this work degenerate to the FBL results for Rayleigh channels when the number of scatterers approaches infinity, the IBL results for Rayleigh-product channels when the blocklength approaches infinity, and the IBL results for Rayleigh channels when both the number of scatterers and the blocklength go to infinity. The derived results can be extended to double-scattering channels and IRS-aided MIMO channels with Rayleigh fading [48] by the approach in Appendix C with the aid of the computation results in [31]. Considering challenges in CSI acquisition for large-scale MIMO systems, it is also meaningful to extend the results to the case with imperfect CSI [49].
@@ Appendix B Mathematical Tools and Useful Results
1. Nash-Poincaré Inequality [19, Eq. (18)][35, Proposition 2.5]. Denote $\mathbb{x}=[x_{1},...,x_{N}]^{T}$ as a complex Gaussian random vector satisfying $\mathbb{E}[\mathbb{x}]=\mathbb{0}$, $\mathbb{E}[\mathbb{x}\mathbb{x}^{T}]=\mathbb{0}$, and $\mathbb{E}[\mathbb{x}\mathbb{x}^{H}]=\mathbb{\Omega}$. $f=f(\mathbb{x},\mathbb{x}^{*})$ is a $\mathbb{C}^{1}$ complex function such that both itself and its derivatives are polynomially bounded. Then, the variance of $f$ satisfies the following inequality,
@@
$\mathbb{E}[x_{i}f(\mathbb{x},\mathbb{x}^{*})]=\sum_{m=1}^{N}[\mathbb{\Omega}]_{i,m}\mathbb{E}\Bigl{[}\frac{\partial f(\mathbb{x},\mathbb{x}^{*})}{\partial x_{m}^{*}}\Bigr{]}.$ (55)
@@
$\mathbb{E}[x_{i}f(\mathbb{x},\mathbb{x}^{*})]=\mathbb{E}\Bigl{[}\frac{\partial f(\mathbb{x},\mathbb{x}^{*})}{\partial x_{m}^{*}}\Bigr{]}.$ (56)
@@
Given Assumption A and defining ${\mathbb{Q}}(z)=(z{\mathbb{I}}_{N}+{\mathbb{H}}{\mathbb{H}}^{H})^{-1}$, for finite $x,z>0$, $m\geq 0$ and matrix ${\mathbb{B}}$, we have
$\mathrm{Var}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{B}}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m})}{M}\Bigr{)}$ $\leq\frac{K_{m}\operatorname{Tr}({\mathbb{B}}{\mathbb{B}}^{H})(1+2mz^{-2}+2x^{-2}]}{M^{3}z^{2m}x^{2}},$ (57)
$\mathrm{Var}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}{\mathbb{H}}{\mathbb{B}}{\mathbb{H}}^{H})}{M}\Bigr{)}$ $\leq\frac{K_{m}\operatorname{Tr}({\mathbb{B}}{\mathbb{B}}^{H})(3+2x^{-2}+2mz^{-2}]}{M^{3}z^{2m}x^{2}},$ (58)
$\mathrm{Var}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{Q}}(x){\mathbb{Q}}(z)^{m}{\mathbb{H}}{\mathbb{B}}{\mathbb{H}}^{H})}{M}\Bigr{)}$ $\leq\frac{K_{m}\operatorname{Tr}({\mathbb{B}}{\mathbb{B}}^{H})(2+2x^{-2}+2mz^{-2})}{M^{3}z^{2m}x^{2}},$ (59)
$\mathrm{Var}\Bigl{(}\frac{\operatorname{Tr}({\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}({\mathbb{Q}}(z){\mathbb{H}}{\mathbb{B}}{\mathbb{H}}^{H})^{m})}{M}\Bigr{)}$ $\leq\frac{K_{m}\operatorname{Tr}(({\mathbb{B}}{\mathbb{B}}^{H})^{m})(2m+1+2x^{-2}+2mz^{-2})}{M^{3}x^{2}z^{2m}},$ (60)
$\mathrm{Var}\Bigl{(}\frac{\operatorname{Tr}(({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})^{m}{\mathbb{Q}}(x){\mathbb{H}}{\mathbb{B}}{\mathbb{H}}^{H})}{M}\Bigr{)}$ $\leq\frac{K_{m}\operatorname{Tr}({\mathbb{B}}{\mathbb{B}}^{H})(m+2+2x^{-2}+2mz^{-2})}{M^{3}x^{2}z^{2m}}.$ (61)
@@
$\mathrm{Var}\Bigl{(}\frac{\operatorname{Tr}{\mathbb{B}}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}}{M}\Bigr{)}\leq\frac{1}{M^{2}L}\mathbb{E}\Bigl{[}\sum_{i,j}|[{\mathbb{Y}}{\mathbb{H}}^{H}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}{\mathbb{B}}{\mathbb{Q}}(x)]_{j,i}|^{2}+|[{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}{\mathbb{B}}{\mathbb{Q}}(x){\mathbb{H}}{\mathbb{Y}}^{H}]_{i,j}|^{2}$ (62)
@@
$+|[{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}{\mathbb{B}}{\mathbb{Q}}(x)]_{j,i}|^{2}+|[{\mathbb{Q}}(z)^{m}{\mathbb{B}}{\mathbb{Q}}(x){\mathbb{Z}}]_{i,j}|^{2}\Bigr{]}$
$+\frac{1}{M^{3}}\mathbb{E}\Bigl{[}\sum_{i,j}|[{\mathbb{H}}^{H}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}{\mathbb{B}}{\mathbb{Q}}(x){\mathbb{Z}}]_{j,i}|^{2}+|[{\mathbb{Z}}^{H}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m}{\mathbb{B}}{\mathbb{Q}}(x){\mathbb{H}}]_{i,j}|^{2}$
$+\sum_{k=1}^{m}|[{\mathbb{H}}^{H}{\mathbb{Q}}^{k}(z){\mathbb{B}}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m-k+1}{\mathbb{Z}}]_{j,i}|^{2}+|[{\mathbb{Z}}^{H}{\mathbb{Q}}^{k}(z){\mathbb{B}}{\mathbb{Q}}(x){\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z)^{m-k+1}{\mathbb{H}}]_{i,j}|^{2}\Bigr{]}$
$\leq\frac{1}{M^{3}}[2K_{1}z^{-2m}x^{-4}\operatorname{Tr}({\mathbb{B}}{\mathbb{B}}^{H})+2K_{1}mz^{-2(m+1)}x^{-2}\operatorname{Tr}({\mathbb{B}}{\mathbb{B}}^{H})+2\operatorname{Tr}({\mathbb{B}}{\mathbb{B}}^{H})x^{-2}z^{-2m}$
$+2K_{2}z^{-2m}x^{-4}\operatorname{Tr}({\mathbb{B}}{\mathbb{B}}^{H})+2K_{2}mz^{-2(m+1)}x^{-2}\operatorname{Tr}({\mathbb{B}}{\mathbb{B}}^{H})]$
$\leq\frac{K_{m}\operatorname{Tr}(({\mathbb{B}}{\mathbb{B}}^{H})^{m})(2m+1+2x^{-2}+2mz^{-2})}{M^{3}x^{2}z^{2m}},$
@@
$\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]}{M}=\frac{\omega\delta_{z}}{\delta(1+\delta_{z}\overline{\omega}\overline{\omega}_{z})\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)},$ (67)
$\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})]}{M}=\frac{\overline{\omega}\overline{\omega}_{z}\omega_{z}\delta_{z}}{\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)},$ (68)
@@
$=\frac{\delta\overline{\omega}\omega_{z}\overline{\omega}_{z}}{1+\delta\overline{\omega}\overline{\omega}_{z}}+\frac{M\overline{\omega}\omega_{z}^{2}\overline{\omega}_{z}}{L\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})^{2}\Delta_{\sigma^{2}}(z)}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)},$
@@
$+\frac{M\omega_{z}}{L\delta_{z}(1+\delta\overline{\omega}\overline{\omega}_{z})}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}{\mathbb{Q}}(z))]}{M}++{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)},$
@@
$\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})]}{M}$ (73)
$=\frac{1}{M}\sum_{i,j}\mathbb{E}[Y_{j,i}^{*}[{\mathbb{Z}}^{H}{\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}]_{j,i}]=\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}-\mathbb{E}\Bigl{[}\frac{\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})}{M}\frac{\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})}{M}\Bigr{]}$
$-\mathbb{E}\Bigl{[}\frac{\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})}{M}\frac{\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})}{M}\Bigr{]}=\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}-\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})]}{M}$
$-\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})]}{M}+\varepsilon_{H,1}+\varepsilon_{H,2},$
where $\varepsilon_{H,1}=-\frac{1}{M^{2}}\mathrm{Cov}(\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H}),\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}))$ and $\varepsilon_{H,2}=-\frac{1}{M^{2}}\mathrm{Cov}(\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}),\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}))$. By using the Cauchy-Schwarz inequality and the variance control in (57) and (59) of Lemma 2, $|\varepsilon_{H,1}|$ and $|\varepsilon_{H,2}|$ can be bounded by
$|\varepsilon_{H,1}|\leq\frac{1}{M^{2}}\mathrm{Var}^{\frac{1}{2}}(\operatorname{Tr}{\mathbb{Q}}(z){\mathbb{Z}}{\mathbb{Z}}^{H})\mathrm{Var}^{\frac{1}{2}}(\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}))={\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(z^{-1})}{M^{2}z^{2}}\Bigr{)},$ (74)
$|\varepsilon_{H,2}|\leq\frac{1}{M^{2}}\mathrm{Var}^{\frac{1}{2}}(\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H}))\mathrm{Var}^{\frac{1}{2}}(\operatorname{Tr}({\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H}))={\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(z^{-1})}{M^{2}z}\Bigr{)},$
@@
$\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})]}{M}=(1-\omega\overline{\omega})\frac{\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}-\frac{\omega_{z}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{H}}{\mathbb{H}}^{H})]}{M}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)}$ (75)
=$\overset{(b)}{=}\frac{\overline{\omega}\overline{\omega}_{z}\mathbb{E}[\operatorname{Tr}({\mathbb{Q}}(z){\mathbb{Q}}{\mathbb{Z}}{\mathbb{Z}}^{H})]}{M}+{\mathcal{O}}\Bigl{(}\frac{\mathcal{P}(\frac{1}{z})}{zM^{2}}\Bigr{)},$